# DiffoRA: Enabling Parameter-Efficient LLM Fine-Tuning via Differential Low-Rank Matrix Adaptation

### v1

@@ 4.1 Continuous relaxation
Figure 1: The framework of DiffoRA, contains two stages. In stage one (left part), the initialized adaptive matrix is continuously relaxed, i.e., C.R. in the figure. The DAM is then differentiable and can be updated. In stage two (right part), the obtained DAM is discretized to binary. All the modules corresponding to entry one will be fine-tuned using decomposition matrices.
The first stage of DiffoRA is illustrated in the left part of Fig. 1. In contrast with the existing work, we point out that it is unnecessary to allocate the rank for every module. Instead, we can construct a selective matrix called DAM module-wisely that determines which module needs to be fine-tuned and which need not. The final output DAM $\bm{\Gamma}$ should be a binary matrix, in which the “ones” (resp. “zeros”) indicate that the corresponding entries will (resp. will not) be fine-tuned in the following procedures. However, directly generating a binary matrix is non-trivial and lacks foundation, i.e., it is an NP-hard problem. To this end, in our design of DiffoRA, we first relax the range of the elements in $\bm{\Gamma}\in\{0,1\}^{L\times N}$ to $\bm{\bar{\Gamma}}\in[0,1]^{L\times N}$ continuously and view all the row vectors in $\bar{\bm{\Gamma}}$ as the hyperparameters, which can be differentiated and updated.
More specifically, for the collection of the modules in the $i$-th layer, we utilize the row vector $\bar{\bm{\gamma}}^{i}\in[0,1]^{1\times N}$ of $\bm{\bar{\Gamma}}$ as learnable hyperparameters in continuous space. The forward pass of the relaxed LoRA-based fine-tuning formula can then be defined as:
$h_{j}^{i}=\bm{W}_{j}^{i}\bm{x}+\bar{\gamma}_{j}^{i}\Delta\bm{W}_{j}^{i}\bm{x}=\bm{W}_{j}^{i}\bm{x}+\bar{\gamma}_{j}^{i}\bm{B}_{j}^{i}\bm{A}_{j}^{i}\bm{x},$ (10)
where $h_{j}^{i}$ denotes the $j$-th item of the hidden nodes in the $i$-th layer, and each $\bar{\gamma}_{j}^{i}$ in $\bar{\bm{\gamma}}^{i}$ satisfies
$\bar{\gamma}_{j}^{i}=\frac{\exp(\gamma_{j}^{i})}{\sum_{j\in[N]}\exp(\gamma_{j}^{i})},$ (11)
$\bm{W}_{j}^{i}\in\mathbb{R}^{d\times k}$ is the pre-trained weight matrix of module $m_{j}^{i}$. $\bm{B}_{j}^{i}$ and $\bm{A}_{j}^{i}$ are $j$-th low-rank decomposition matrices in the $i$-th layer.
Intuitively, $\bm{\bar{\Gamma}}$ can represent the necessities of fine-tuning this module using $\mathbf{A}$ and $\mathbf{B}$. When this weight tends to $1$, it indicates that fine-tuning this module is necessary and might lead to better performance, and vice versa. By utilizing the continuous relaxation on the differential matrix, we are able to generate each $\bar{\bm{\gamma}}^{i}$ via a few rounds of training (e.g., five rounds, determined by the datasets). We denote $\mathcal{L}_{train}$ and $\mathcal{L}_{valid}$ as the training and validation loss, both of which are functions of $\Delta\bm{W}$ and $\bar{\bm{\Gamma}}$ such that
$\mathcal{L}_{\cdot}:=\mathcal{L}_{\cdot}(\Delta\bm{W},\bar{\bm{\Gamma}})$ (12)
Our goal is to find the best hyperparameters $\bar{\bm{\Gamma}}$ to minimize the validation loss, which is equivalent to bi-level optimization problems as follows:
To solve the optimizations in Eq. 13, we use the gradient descent algorithm to update the parameters $\bm{A}$, $\bm{B}$ and $\bar{\bm{\Gamma}}$. Specifically, we randomly extract part of the data in the training set and divide it into two parts, i.e., the training data and validation data, which are used to update $\Delta\bm{W}$ and $\bar{\bm{\Gamma}}$ alternately. We present the detailed training algorithms in Algorithm 1, lines 1 to 8. Note that during this step, the remaining network parameters are kept frozen as pre-trained.
4.2 Discretization and fine-tuning
Having the relaxed matrix $\bar{\bm{\Gamma}}$ in the previous stage, we perform the discretization to obtain the binary DAM $\bm{\Gamma}$. An intuitive representation of this procedure is shown in the right part of Fig. 1. More concretely, for each item in $\bar{\bm{\Gamma}}$, we update the top $K:=\lfloor\rho\cdot N\rfloor$ largest entries with 1 and set the remaining values as 0, such that
$\gamma^{i}_{j}=\begin{cases}1&\mbox{if $\bar{\gamma}^{i}_{j}\geq\delta^{i}$}\\ 0&\mbox{otherwise}\end{cases}$ (14)
where $\delta^{i}$ is the value of the $K$-th largest entry in $\bar{\bm{\gamma}}^{i}$ and $\rho$ is the selecting ratio. In our design, $\rho$ is a hyperparameter which is set as 0.5 for fair comparison with the baselines. The discretization is described in Algorithm 1, lines 10 and 11. The final step is to fine-tune the model equipped with the binarized DAM and low-rank decomposition matrices. During the fine-tuning, only the modules with weight one (e.g., the selected module) will be fine-tuned, while the others are kept unchanged in the downstream tasks.
4.3 Optimization and weight sharing
In the previous section, we obtain the DAM $\bm{\Gamma}$ using continuous relaxation and discretization. This approach can significantly improve the model performance after fine-tuning when the discrepancy of the weights in $\bm{\bar{\Gamma}}$ is distinct. However, when the entries in the relaxed $\bm{\bar{\Gamma}}$ display a uniform distribution, the method in the previous descriptions can be further optimized. More concretely, as shown in Fig. 2, we take the continuous DAM on CoLA and MRPC [31] as examples. We visualize matrix $\bar{\bm{\Gamma}}$ on these two datasets with DeBERTaV3-base [9] as the backbone, respectively, in which each column is a layer with six trainable modules. It can be seen that the weight distribution on CoLA is relatively distinct, i.e., in the majority of the layers, $W_{O}$ and $W_{Q}$ obtain significantly higher weights so that they are more suitable for the following fine-tuning. In contrast, the entries of the $\bm{\bar{\Gamma}}$ on MRPC share a similar and uniform distribution. For instance, all the candidate modules obtain weights around 0.15 from layers 1 to 8. Under this circumstance, if we select the top-$K$ largest entries, the modules that correspond to the “0” entries after discretization also demonstrate the same level of importance as those selected by ones. Consequently, directly fixing the modules with zero entries in $\mathbf{\Gamma}$ might lead to performance degradation.
Figure 2: The module weights of the DeBERTaV3-base model in the QNLI and RTE datasets. In the figure, $W_{Q}$, $W_{K}$, $W_{V}$, $W_{I}$, $W_{O}$ and $W_{D}$ correspond to the $\mathsf{query\_proj}$, $\mathsf{key\_proj}$, $\mathsf{value\_proj}$, $\mathsf{intermediate.dense}$, $\mathsf{output.dence}$, and $\mathsf{attention.output.dense}$ modules in the pre-trained model respectively.
To address this discretization discrepancy issue, we adopt the weight-sharing strategy to further optimize our method. Specifically, for the modules corresponding to the zero entries after discretization, instead of just freezing them without fine-tuning, we fine-tune them with the same weights as the modules in other layers. In other words, those modules with zero DAM entries share the same model weights in the fine-tuning procedure. For instance, all the modules $\bm{W}_{i}$ in layers 1 to 12 that are not selected will share the same weights and also participate in fine-tuning, where $\bm{W}_{i}\in\{\bm{W}_{Q},\bm{W}_{K},\bm{W}_{V},\bm{W}_{I},\bm{W}_{O},\bm{W}_{D}\}$. This optimization can enhance the performance of DiffoRA without introducing a large amount of extra fine-tuning parameters. The model fine-tuning with weight-sharing is shown in Algorithm 1, lines 12 to 15.
5 Experiments
5.1 Configurations
Hardware. DiffoRA is fully implemented in Python programming language. We evaluate our method on a Desktop Core i7-12700F CPU and GeForce RTX 3090.
Datasets and pre-trained models. We utilize two types of benchmarks: i) General Natural Language Understanding (GLUE) [30], including MNLI, SST-2, CoLA, QQP, QNLI, RTE, MRPC, and STS-B; and ii) Question Answering, including SQuADv1.1 [26] and SQuADv2.0 [27]. We use the DeBERTaV3-base [9] as the backbone model in the main text, and the results on GLUE using RoBERTa-base [20] are presented in Appendix D.
Counterpart comparisons. We use the following methods as baselines. (i) Full FT uses all parameters for fine-tuning; (ii) BitFit [35] is a sparse-fine-tuning method for pre-trained models that updates only a small subset of the bias terms; (iii) Houlsby adapter [11] adds a few trainable modules inserted between layers of a pre-trained model, allowing for task-specific tuning without altering the entire model; (iv) Pfeiffer adapter [24] combines multiple task-specific adapters by linearly blending their outputs; (v) LoRA [12] reduces the number of trainable parameters by applying low-rank matrix decomposition to weight updates in pre-trained models; (vi) AdaLoRA [36] adapts LoRA by dynamically adjusting the rank of low-rank updates during training, optimizing parameter efficiency while maintaining model performance across various tasks. We select the baseline methods with comparable parameter amounts and open-sourced codes for fair comparisons.
Implementation details. We set the module retention ratio $\rho$ to 50%, the LoRA rank to 4, and the $\alpha$ of LoRA to 16. All the results are the average values under three random seeds. See Appendix C for more detailed settings.
5.2 Natural language understanding
We evaluate the performance of the fine-tuned model on the GLUE benchmark [31] using various approaches. We use DeBERTaV3-base as the pre-trained model, which contains 183 million parameters. During this task, we set the rank of the low-rank decomposition matrices to 4 and fine-tuned 50% of the modules in each layer of the pre-trained model. We adopt all eight datasets in GLUE and the concrete fine-tuning model architectures are determined by the specific tasks. We summarize and present the results in Tab. 1. Overall, our DiffoRA achieves the highest model accuracy among the baselines on all the datasets. More concretely, our method achieves 0.81% higher fine-tuning accuracy than the state-of-the-art method AdaLoRA on the CoLA dataset. DiffoRA exhibits consistently better results than the existing methods under the same level of parameter amounts. Moreover, our method also outperforms the baseline methods with a larger parameter size, i.e., up to 3.84% higher accuracy than PAdapter with twice the parameters involved. Compared to the full fine-tuning strategy, DiffoRA obtains a better performance, which also demonstrates the effectiveness of our approach in the GLUE benchmark.
Method #Params MNLI SST-2 CoLA QQP QNLI RTE MRPC STS-B All
m/mm Acc Mcc Acc/F1 Acc Acc Acc Corr Avg.
AdaLoRA 0.32M 90.09/90.41 95.80 70.04 91.78/89.16 94.49 87.36 90.44 91.63 88.81
DiffoRA 0.35M 90.49/90.49 96.09 70.85 91.79/89.12 94.52 87.96 90.79 91.75 89.11
Table 1: The results of the fine-tuned DeBERTaV3-base model on the GLUE dataset are presented, with the best results highlighted in bold and the second-best results underlined. Our DiffoRA achieved the best results on average.
5.3 Question answering
For the question-answering task, we use DeBERTaV3-base as the pre-trained model and adopt two datasets (i.e., SQuADv1.1 and SQuADv2.0) under different amounts of parameters to fine-tune the model. In order to keep the number of parameters close to the baseline, we choose the rank of the low-rank decomposition matrices from $\{1,2,5,10\}$ in the SQuADv1.1 dataset and $\{2,4,8,15\}$ in the SQuADv2.0 dataset. We use Exact Match (EM) and F1 as evaluation indicators and summarize the results in Tab. 2. Similar to the GLUE benchmark, our DiffoRA also demonstrates consistently better results than the baseline on both SQuAD datasets. Specifically, on SQuADv1.1, DiffoRA obtains 0.4% to 0.5% higher EM and 0.1% to 0.2% higher F1 than the best baseline AdaLoRA. Compared to the full fine-tuning method results, our DiffoRA also achieves higher accuracy even with 0.08% parameters fine-tuned. Furthermore, on the SQuADv2.0 dataset, our method is around 0.2% higher than the best baseline on EM and F1. DiffoRA is consistently better than the fully fine-tuning strategy (i.e., 85.4% EM and 88.4% F1) and original LoRA, which demonstrates the effectiveness of our scheme.
In all, the experimental results in this section demonstrate that DiffoRA works consistently better than the baseline methods on all benchmarks and datasets, which conforms to our theoretical analysis.
Method SQuADv1.1 SQuADv2.0
Full FT 86.0/92.7 85.4/88.4
#Params 0.08% 0.16% 0.32% 0.65% 0.08% 0.16% 0.32% 0.65%
6 Analysis
DiffoRA v.s. Random select strategy. We first analyze the effectiveness of DiffoRA by comparing its performance with that of the random selection method. Specifically, we randomly select three modules in each layer as the fine-tuning modules, and the remaining modules are processed with the weight-sharing strategy. The results are shown in Tab. 3. We choose ranks from $\{1,2,4\}$ for fine-tuning in the STS-B and SQuADv1.1 datasets, respectively. It can be demonstrated that compared to random sampling, DiffoRA achieves significant performance improvements and effectively identifies important modules, e.g., DiffoRA is 0.2% to 0.6% higher than the random select strategy.
Method Rank #Params STS-B SQuADv1.1
Random 1 0.09 91.21 87.23
DiffoRA 1 0.08 91.65 87.37
Random 2 0.19 91.38 87.55
DiffoRA 2 0.16 91.52 87.74
Random 4 0.37 91.11 87.82
DiffoRA 4 0.34 91.75 88.04
Table 3: Random Selection v.s. DiffoRA on two datasets.
Sample rate. We select sample rates from {0.2, 0.4, 0.5, 0.7, 0.9} corresponding to the most important Top-$\{1,2,3,4,5\}$ modules, respectively. To explore the relationship between the performance and sample rate, we conduct experiments on three datasets: STS-B, RTE, and MRPC. The results are summarized in Tab. 4. The results show that when the sampling rate is around 0.5, the performance of the fine-tuned model achieves state-of-the-art.
Sample Rate K STS-B RTE SQuADv1.1
0.2 1 90.87 85.07 87.09
0.4 2 91.66 85.65 87.43
0.5 3 91.75 87.96 88.12
0.7 4 91.41 87.04 88.19
0.9 5 91.39 86.34 88.18
Table 4: DiffoRA across three datasets at different sample rates.
Weight sharing. We further investigate the effects of the weight-sharing strategy on our DiffoRA. Specifically, we consider two scenarios: (i) weight sharing, and (ii) weight sharing combined with a selection matrix. Experiments are conducted on the STS-B and SQuADv1.1 datasets, and the results are summarized in Tab. 5. The table shows that the combination of module selection and weight sharing can lead to better results. For instance, our DiffoRA with the weight sharing achieves 0.11% to 2.9% higher accuracy than the method without this strategy.
Weight Share STS-B RTE MRPC
91.75 87.96 90.79
91.66 85.06 89.58
Table 5: Comparison of DiffoRA w/o weight sharing.
7 Conclusion
We propose a new PEFT method called DiffoRA, which enables efficient and adaptive LLM fine-tuning based on LoRA. Instead of adjusting every interior rank, we argue that adopting LoRA module-wisely is sufficient. To achieve this, we construct a DAM to select the modules that are most suitable and essential to fine-tune. We theoretically analyze how the DAM impacts the convergence rate and generalization capability. Furthermore, we adopt continuous relaxation and discretization to establish DAM. To alleviate the issue of discretization discrepancy, we utilize the weight-sharing strategy for optimization. The experimental results demonstrate that our DiffoRA works consistently better than the baselines across all benchmarks.
Supplementary Material
Appendix A Detailed Proofs
A.1 Proof of Theorem 2
We first introduce the Wely inequality as follows.
Lemma 1 (Weyl inequality [10]).
We first present and proof the following lemma.
Lemma 2.
If $\bm{x}_{i}\nparallel\bm{x}_{j},\forall i\neq j$, we have $\bm{H}^{\infty}_{w_{0}}\succ 0$.
Proof.
By the Lemma 3.4 of [4], there exists $\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})$, such that when $m$ is sufficiently large, $\|\bm{w}-\bm{w}_{0}\|$ is sufficiently small. Then according to the proof of Theorem 3.1 of [4], we get $\lambda_{\min}(\bm{H}_{w_{0}}^{\infty})>0$.
Then we provide the proof of Theorem 2.
Theorem 2. Suppose $f$ is an NN with a single hidden layer and ReLU activation function. Assume $\bm{X}\in\mathbb{R}^{d\times n}$, $\mathbf{w}(0)\sim N(\mathbf{0},\mathbf{I})$, hidden nodes $m=\Omega\left(\frac{n^{6}d^{2}}{(\lambda_{0}^{\Gamma})^{4}\delta^{3}}\right)$, and $\bm{I}^{\bm{\Gamma w}}-\bm{I}^{\bm{w}}\succeq 0$, then the following formula holds with probability at least $1-\delta$ over the initialization
$\leq$ $\exp(-\lambda_{0}^{\Gamma}t)\|f(\mathbf{W}(0),\mathbf{a},\mathbf{X};\bm{\Gamma},\bm{W}_{0})-\bm{y}\|_{2}^{2}$
where $\lambda_{0}^{\Gamma}\geq\lambda_{0}$.
Proof.
We denote $\bm{I}^{-}:=\bm{I}^{\bm{\Gamma w}}-\bm{I}^{\bm{w}}\succeq 0$. Then we have the following inequalities:
$\lambda_{\min}(\bm{H}^{\infty}_{\Gamma,w_{0}})$ $=\lambda_{\min}(\bm{H}^{\infty}_{w_{0}}+\bm{H}^{\infty}_{\Gamma,w_{0}}-\bm{H}_{w_{0}}^{\infty})$ (18)
$\geq\lambda_{\min}(\bm{H}_{w_{0}}^{\infty})+\lambda_{\min}(\bm{H}^{\infty}_{\Gamma,w_{0}}-\bm{H}_{w_{0}}^{\infty})$
From the definitions of $\bm{H}^{\infty}_{\Gamma,w_{0}}$ and $\bm{H}^{\infty}_{w_{0}}$ we have
$\bm{H}^{\infty}_{\Gamma,w_{0}}-\bm{H}_{w_{0}}^{\infty}=\bm{X}^{T}\bm{X}\odot\bm{I}^{-}$ (19)
Since $\bm{X}^{T}\bm{X}$ and $\bm{I}^{-}$ are both positive definite/semi-positive definite matrices, their Hadamard product is also a positive definite/semi-positive definite matrix [22], i.e., $\bm{H}^{\infty}_{\Gamma,w_{0}}-\bm{H}^{\infty}_{w_{0}}\succeq 0$. Therefore, we have
$\lambda_{\min}(\bm{H}_{\Gamma,w_{0}}^{\infty})\geq\lambda_{\min}(\bm{H}_{w_{0}}^{\infty})$ (20)
Finally, according to Theorem 1, we have
$\leq$ $\exp(-\lambda_{0}^{\Gamma}t)\|f(\mathbf{W}(0),\mathbf{a},\mathbf{X};\bm{\Gamma},\bm{W}_{0})-\bm{y}\|_{2}^{2}$
where $\lambda_{0}^{\Gamma}\geq\lambda_{0}$. ∎
Datasets learning rate batch size epochs share r r K $\alpha$ LoRA drop warm-up early stop
MNLI 3e-4 196 7 2 4 3 16 0.25 4000 -
RTE 7e-4 32 50 2 4 3 16 0.2 900 -
QNLI 6e-4 64 5 1 6 3 16 0.35 3000 -
MRPC 1e-3 16 14 2 4 3 16 0 900 -
QQP 8e-4 64 5 2 4 3 16 0 2000 -
SST-2 1e-4 32 5 2 4 3 16 0 3000 -
CoLA 3e-4 16 6 2 4 3 16 0.15 900 5
STS-B 7e-4 16 7 2 4 3 16 0 900 -
Table 6: Training settings for GLUE benchmarks.
Settings SQuADv1.1 SQuADv2.0
#Params 0.08% 0.16% 0.32% 0.65% 0.08% 0.16% 0.32% 0.65%
r 1 2 5 10 2 4 8 15
train epochs 3 3
learning rate 5e-4 7e-4
warm-up 2000 4000
share r 1 2
K 3 3
$\alpha$ 16 16
LoRA drop 16 16
Table 7: Training settings for SQuAD benchmarks.
A.2 Proof of Theorem 3
Theorem 3. For an over-parameterized neural network with the loss on the testing set as $\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})$. Let $\bm{y}=(y_{1},...,y_{N})^{T}$, and $\eta=\kappa C_{1}\sqrt{\bm{y}^{T}(\bm{H}_{\Gamma,w_{0}}^{\infty})^{-1}\bm{y}}/(m\sqrt{N})$ for some small enough absolute constant $\kappa$, where $\eta$ denotes the step of SGD. Under the assumption of Theorem 2, for any $\delta\in(0,e^{-1}]$, there exists $m^{\ast}(\delta,N,\lambda_{0}^{\Gamma})$, such that if $m\geq m^{*}$, then with probability at least $1-\delta$, we have
[…]

### v2

@@ Appendix E Algorithm pseudocode
Appendix F Results on RoBERTa-base
Finally, to further prove the effectiveness of our proposed DiffoRA, we adopt RoBERTa-base as the backbone model which contains 125 million parameters. We fine-tune the pre-trained model on eight tasks of the GLUE benchmark. We set the rank of the low-rank decomposition matrices to $4\sim 12$ and fine-tuned $20\%\sim 70\%$ of the modules in each layer of the pre-trained model. The experimental results are summarized in Table 8. It can be shown that our DiffoRA achieves the best results in almost all datasets. For instance, our approach outperforms the state-of-the-art method PAdapter on the CoLA dataset by 0.98%. Although DiffoRA achieves the second-best result on a few tasks, we achieve the highest average accuracy among all the baselines, i.e., 86.81% on average. The experimental results demonstrate the effectiveness of our method.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our contributions and scope can be accurately reflected in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations can be found in Section 7.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide the full set of assumptions and a complete proof in the Appendix A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provided a README file in our code for the reviewers to check.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provided an anonymous URL link for our codes in Section 5.1. All the datasets we used are public.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have described our experimental settings in Section 5.1. For additional experimental settings on various benchmarks, we provided the details in Appendix D.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We have clearly stated the factors of variability that the error bars are capturing. We have run our experiments multiple times with different seeds and provided the mean and standard deviation values in Section 5.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We have provided the hardware configurations in Appendix D.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We confirm that we preserve the anonymity and conform to the NeurIPS Codes of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This work is fundamental research in computer science and has no societal impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no risks above.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All the codes and datasets used in this paper are open-sourced. We have properly cited these resources in our paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We have provided an anonymous URL link for our codes.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
[…]